# Compression for Adaptive Edge LLMs

### v2

@@ C.1 Ablation study on calibration sets
Table 14 presents the performance of Llama-3.1-8B under different combinations of calibration sets at a fixed 25% pruning rate. Following our setting, we report the average accuracy of five zero-shot downstream tasks. For each configuration, all hyperparameters and the number of calibration samples strictly follow the settings in Table 11 of our manuscript, ensuring a controlled and consistent comparison. Using the Alpaca dataset [taori2023stanford] for all stages results in the best average accuracy. We follow MoDeGPT and use WikiText2 as the calibration set for pruning-ratio allocation to ensure a fair comparison with prior work in all experiments.
@@ Appendix D Pareto-front Analysis
Figure 5 and 6 illustrate the Pareto-frontier trade-offs between accuracy and latency across a diverse set of Transformer, hybrid, and state-space models on A6000 and Nano 8G, respectively. For the A6000, the latency is profiled with 1024 prefilling tokens and 1024 generated tokens. On Orin Nano, we report that the latency of the request, prefilled with 512 tokens, generates 512 new tokens in seconds (sec.). Each subplot groups models by: (a) pure Transformers, (b) hybrid and SSM-based models, and (c) the union of both. We compare UniQL (W4A16, starred markers), GPTQ [frantar2023gptq] (W4A16, circular markers), and FP16 baselines (squares), with circle sizes indicating memory consumption on A6000. For the Nano 8G, we use HQQ [badri2023hqq] that is supported in the TorchAO framework [torchao] (TAO-HQQ) as our baseline. Across all architectures, UniQL consistently yields better latency–accuracy trade-offs under the same memory budget, especially in the 2–4GB regime critical for edge deployment. For example, UniQL significantly improves latency for Qwen-2.5-7B and Mamba-2-8B while maintaining accuracy close to the FP16 baseline. Notably, UniQL achieves Pareto-dominant points for models like Llama-3.1-8B, outperforming GPTQ and TAO-HQQ in both latency and accuracy. Our analysis underscores UniQL’s advantage for high-performance LLM inference under tight latency and memory constraints.
@@ Appendix E Energy Profiling
Method
W-bit
Prun.
p%
Qwen-2.5-7B Mamba2-8B
@@ Appendix F Layer-wise Pruning Rates
We adopt the approach from [lin2025modegpt, men2024shortgpt] to determine layer-wise pruning rates $r_{l}$ using Block Influence (BI) scores for specified global pruning rates. The BI score is given by $s=1-\mathbb{E}\frac{\mathbf{x}_{l}^{\top}\mathbf{y}_{l}}{\|\mathbf{x}_{l}\|_{2}\|\mathbf{y}_{l}\|_{2}}$, with $x_{l}$ and $y_{l}$ as the input and output of the block (e.g., Transformer or Mamba block) at layer $l$. Using the closed-form solution from lin2025modegpt, we smooth the layer-wise pruning rate allocations to obtain $P=[r^{P}_{1},r^{P}_{2},\ldots,r^{P}_{L}]$, such that $P=LP_{\text{avg}}\times\text{Softmax}(-\mathbf{s}/\varepsilon)$ where $s_{i}$ represents the importance score of layer $i$, and $P_{\text{avg}}$ denotes the target global sparsity. In our experiments, $\varepsilon$ is set to $0.1$, and we present the pruning rates for all models in Figure 8. Self-attention layers in hybrid models have low pruning rates, indicated by high BI scores in the Figure. In Bamba-9B-v2, layers $9$, $18$, and $27$ are self-attention layers with lower pruning rates than nearby layers. Similarly, Nemotron-H-8B shows this pattern in layers $7$, $18$, $29$, and $40$. Pruning self-attention layers in hybrid models leads to significant accuracy drops.
@@ G.1 Calibration sets
Table 17 lists the calibration set, number of samples, and the sequence length we use in our experiments. We collect BI scores and assign layer-specific pruning rates using 128 samples with a sequence length of 2048 from wikitext-2 [merity2017pointer]. Various global pruning rates, such as $P_{15}$, $P_{25}$, and $P_{35}$, are computed using the same setting. With 128 samples and a sequence length of 2048, we compute channel correlations from the Alpaca dataset [taori2023stanford], as detailed in [lin2025modegpt]. Our masked LoRA fine-tuning is conducted on the Alpaca dataset with a sequence length of 256 to reduce training memory usage. Lastly, we calibrate post-training quantization on the wikitext-2 dataset using 256 samples with a sequence length of 2048.
@@ G.2 Hyper-parameters of Masked LoRA fine-tuning
We list the hyper-parameters for our masked LoRA fine-tuning in Table 18. We follow the prior work to perform instruction tuning on the Alpaca dataset [taori2023stanford] for five epochs. Specifically, we adopt a relatively small sequence length of 256 to reduce training cost, and set the LoRA rank $r=8$ with scaling factor $\alpha=16$. A warmup of 100 steps is applied with the AdamW optimizer, and micro-batching is used to accommodate GPU memory limits. All models we experiment with are using the same hyperparameters, and we do not tune the parameters for the model and experiments.

### v3

@@ C.1 Ablation study on calibration sets
Table 14 presents the performance of Llama-3.1-8B under different combinations of calibration sets at a fixed 25% pruning rate. Following our setting, we report the average accuracy of five zero-shot downstream tasks. For each configuration, all hyperparameters and the number of calibration samples strictly follow the settings in Table 11 of our manuscript, ensuring a controlled and consistent comparison. Using the Alpaca dataset [Taori et al., 2023] for all stages results in the best average accuracy. We follow MoDeGPT and use WikiText2 as the calibration set for pruning-ratio allocation to ensure a fair comparison with prior work in all experiments.
@@ Appendix D Pareto-front Analysis
Figure 5 and 6 illustrate the Pareto-frontier trade-offs between accuracy and latency across a diverse set of Transformer, hybrid, and state-space models on A6000 and Nano 8G, respectively. For the A6000, the latency is profiled with 1024 prefilling tokens and 1024 generated tokens. On Orin Nano, we report that the latency of the request, prefilled with 512 tokens, generates 512 new tokens in seconds (sec.). Each subplot groups models by: (a) pure Transformers, (b) hybrid and SSM-based models, and (c) the union of both. We compare UniQL (W4A16, starred markers), GPTQ [Frantar et al., 2023] (W4A16, circular markers), and FP16 baselines (squares), with circle sizes indicating memory consumption on A6000. For the Nano 8G, we use HQQ [Badri and Shaji, 2023] that is supported in the TorchAO framework [torchao, 2024] (TAO-HQQ) as our baseline. Across all architectures, UniQL consistently yields better latency–accuracy trade-offs under the same memory budget, especially in the 2–4GB regime critical for edge deployment. For example, UniQL significantly improves latency for Qwen-2.5-7B and Mamba-2-8B while maintaining accuracy close to the FP16 baseline. Notably, UniQL achieves Pareto-dominant points for models like Llama-3.1-8B, outperforming GPTQ and TAO-HQQ in both latency and accuracy. Our analysis underscores UniQL’s advantage for high-performance LLM inference under tight latency and memory constraints.
@@ Appendix E Energy Profiling
Method W-bit Prun. p% Qwen-2.5-7B Mamba2-8B
@@ Appendix F Layer-wise Pruning Rates
We adopt the approach from [Lin et al., 2025, Men et al., 2024] to determine layer-wise pruning rates $r_{l}$ using Block Influence (BI) scores for specified global pruning rates. The BI score is given by $s=1-\mathbb{E}\frac{\mathbf{x}_{l}^{\top}\mathbf{y}_{l}}{\|\mathbf{x}_{l}\|_{2}\|\mathbf{y}_{l}\|_{2}}$, with $x_{l}$ and $y_{l}$ as the input and output of the block (e.g., Transformer or Mamba block) at layer $l$. Using the closed-form solution from Lin et al. [2025], we smooth the layer-wise pruning rate allocations to obtain $P=[r^{P}_{1},r^{P}_{2},\ldots,r^{P}_{L}]$, such that $P=LP_{\text{avg}}\times\text{Softmax}(-\mathbf{s}/\varepsilon)$ where $s_{i}$ represents the importance score of layer $i$, and $P_{\text{avg}}$ denotes the target global sparsity. In our experiments, $\varepsilon$ is set to $0.1$, and we present the pruning rates for all models in Figure 8. Self-attention layers in hybrid models have low pruning rates, indicated by high BI scores in the Figure. In Bamba-9B-v2, layers $9$, $18$, and $27$ are self-attention layers with lower pruning rates than nearby layers. Similarly, Nemotron-H-8B shows this pattern in layers $7$, $18$, $29$, and $40$. Pruning self-attention layers in hybrid models leads to significant accuracy drops.
@@ G.1 Calibration sets
Table 17 lists the calibration set, number of samples, and the sequence length we use in our experiments. We collect BI scores and assign layer-specific pruning rates using 128 samples with a sequence length of 2048 from wikitext-2 [Merity et al., 2017]. Various global pruning rates, such as $P_{15}$, $P_{25}$, and $P_{35}$, are computed using the same setting. With 128 samples and a sequence length of 2048, we compute channel correlations from the Alpaca dataset [Taori et al., 2023], as detailed in [Lin et al., 2025]. Our masked LoRA fine-tuning is conducted on the Alpaca dataset with a sequence length of 256 to reduce training memory usage. Lastly, we calibrate post-training quantization on the wikitext-2 dataset using 256 samples with a sequence length of 2048.
@@ G.2 Hyper-parameters of Masked LoRA fine-tuning
We list the hyper-parameters for our masked LoRA fine-tuning in Table 18. We follow the prior work to perform instruction tuning on the Alpaca dataset [Taori et al., 2023] for five epochs. Specifically, we adopt a relatively small sequence length of 256 to reduce training cost, and set the LoRA rank $r=8$ with scaling factor $\alpha=16$. A warmup of 100 steps is applied with the AdamW optimizer, and micro-batching is used to accommodate GPU memory limits. All models we experiment with are using the same hyperparameters, and we do not tune the parameters for the model and experiments.